3a(4-5x)=4+30x-28
Find the value of a for which the equation is an identity.

Answer by ewatrrr(24785)

 
Hi,
Find the value of a for which the equation is an identity.
3a(4-5x)=4+30x-28
12a-15ax = 30x - 24
12a = -24 |Comparing Like terms
a = -2
& -15ax = 30x
a = -2